Use the five-step problem solving strategy to find the measure of the angle described.
The angle's measure is 60degrees more than that of its complement.
I am having problems with setting it up please help

Found 2 solutions by mananth, ikleyn:
Answer by mananth(16949)

The angle's measure is 60degrees more than that of its complement.
I am having problems with setting it up please let the angle be x
the complement will be 90-x
x= (90-x)+60
x= 90+60-x
2x=120
x=60
the complement will be 30.
mananth@hotmail.com


Answer by ikleyn(53878)

.
Use the five-step problem solving strategy to find the measure of the angle described.
The angle's measure is 60degrees more than that of its complement.
I am having problems with setting it up please help
~~~~~~~~~~~~~~~~~~~~~~~~~~~


        The solution in the post by @mananth is incorrect due to his arithmetic errors.
        I came to bring a correct solution.


Let the angle be x
the complement will be 90-x
x = (90-x) + 60
x = 90 + 60 - x
2x = 150
x = 150/2 = 75 degrees
the complement will be 15 degrees.

Solved correctly.